# Foundation Model Priors Enhance Object Focus in Feature Space for Source-Free Object Detection

### v1

@@ 5.2 Ablation Studies
GSAM [ren2024grounded] (arXiv’24) 46.2 57.2 48.4
ESC-Net [lee2025effective] (CVPR’25) 46.5 57.8 49.3
OVSAM [yuan2024open] (ECCV’24) 46.9 58.8 50.1
@@
Ablation on different components of IRPL. Table 7 presents an ablation study highlighting the impact of different components in IRPL (Equation 5). The peak-adjust transform ($p^{\prime}_{k}$[wang2024epsilon] provides moderate improvement when directly applied to object detection, yielding 55.9 AP. However, better performance is achieved when it is complemented with: i) foreground-background weighting ($\mathbf{w_{\hat{c}}}$), which addresses the inherent imbalance between object and background regions and ii) entropy regularization ($\mathbf{D_{\mathrm{KL}}}$), which mitigates head-class dominance. Together, these components lead to the highest AP of 56.8, demonstrating the effectiveness of IRPL in SFOD.
@@
Table 8: Per-category AP50 results on the target-domain. We report baseline [hao2024simplifying] performance, performance with our method, and the improvement ($\Delta$).
@@ Appendix A.2 Architecture-generality of FALCON-SFOD
To assess the architectural generality of our framework, we integrate FALCON-SFOD with multiple representative detection methods, including both conventional Faster R-CNN-based pipelines [irg, pets] and transformer-based detectors such as Deformable DETR [khanh2024dynamic]. As shown in Table A.1, incorporating our proposed SPAR and IRPL modules consistently improves performance across architectures and domain-shift settings. Notably, the observed gains are obtained without modifying the detector structure or adding inference-time cost, underscoring the plug-and-play nature of our design. These results demonstrate that FALCON-SFOD generalizes well beyond a single detector family, enhancing adaptation robustness across both convolutional and transformer architectures.
@@ Appendix A.3 Reproducibility and Implementation Details
As described in Section A.2, FALCON-SFOD can be seamlessly integrated into a variety of detector architectures, consistently yielding performance gains across settings. For a fair comparison, we preserve all training parameters such as batch size, number of epochs, learning rate, optimization schedule, and data preprocessing pipeline from the corresponding baseline models, applying FALCON-SFOD as an additional adaptation module without altering the underlying training configuration. For the newly introduced hyperparameters, Section A.3.1 demonstrates that a single configuration performs consistently across benchmarks, indicating that FALCON-SFOD exhibits strong robustness and low sensitivity to hyperparameter choices. The complete anonymized implementation of our proposed method is provided as a zip file in the supplementary material.34
@@ A.3.1 Hyperparameter Analysis
Imbalance-aware Noise-Robust Pseudo-Labeling. IRPL addresses the dual challenges of label noise and foreground-background imbalance inherent in source-free object detection. Unlike the standard cross-entropy objective, which can be dominated by mislabeled high-confidence predictions, IRPL utilizes a peak-adjust transform that moderates the student’s logits by adding a large margin $m$ to the most confident class and renormalizing, thereby dampening gradients for easy, clean samples while preserving full corrective signals for uncertain or mislabeled ones. This mechanism provides intrinsic robustness to noisy pseudo-labels. To further stabilize learning under imbalance, IRPL combines this with foreground-background weighting ($w_{\mathrm{fg}},w_{\mathrm{bg}}$) and a mild entropy regularization term weighted by $\gamma$, which discourages over-confident head-class predictions. Following the theoretical insights from [wang2024epsilon], we treat the coefficients $\alpha$ and $\beta$ in
@@
as scaling factors rather than critical hyperparameters. Lemma 3 in [wang2024epsilon] proves that the excess-risk bound of $\alpha L_{\epsilon}+\beta L_{\text{symmetric}}$ is identical to that of $\alpha L_{\epsilon}$ alone, implying that $\alpha$ and $\beta$ affect only the gradient magnitude rather than the theoretical robustness. Hence, we fix them following [wang2024epsilon] instead of tuning.
@@ A.3.2 Results across Random Seeds
Discussion. While existing SFOD works largely report single-run results, Table A.3 provides performance across three random seeds to evaluate robustness and stability. When integrated into Simple-SFOD [hao2024simplifying], our FALCON-SFOD consistently improves performance on the challenging $\rightarrowFoggy Cityscapes$ domain shift. Across seeds 42, 123, and 9999, our approach achieves an average mAP of 47.0 $\pm$ 0.29, demonstrating both a clear improvement over the Simple-SFOD baseline (45.0 mAP) and strong reproducibility. The small variance across runs indicates that FALCON-SFOD’s gains are stable and not sensitive to random initialization.
@@
Simple-SFOD [hao2024simplifying] (ECCV’24) 40.9 48.0 58.9 29.6 51.9 50.2 36.2 44.1 45.0
@@ Appendix A.4 Proofs for the Lemmas and Theorems
Inspired from [wang2024epsilon], we can write
@@ Appendix A.5 Run time and Memory Analysis
Tables A.4 and A.5 report the cost of integrating GSAM [ren2024grounded] into our pipeline. The time overhead is marginal: the offline GSAM pass completes in 1 0501\,050 s ($\sim$17 min), which is only $3.8\%$ of the 28 00028\,000 s required for our model training (and only $5.2\%$ relative to the baseline). When added to the full pipeline, the end-to-end wall-clock increases only slightly (28 08428\,084 s $\rightarrow$ 29 13429\,134 s), well within the typical run-to-run variability of large-scale training. The memory overhead is also short-lived: GSAM peaks at $18$ GB only during its 17-minute preprocessing, while training itself never exceeds $9.6$ GB. Since these stages do not overlap, the entire procedure fits comfortably on a single 2424-$48$ GB GPU without any modification to training. Moreover, the amortization cost is small: GSAM masks are generated once per target split and can be cached for reuse in all subsequent experiments. The extraction step is fully parallel across images, so on a multi-GPU node the elapsed time approaches standard data-loading latency. For completeness, we note that ESC-Net [lee2025effective] and OV-SAM [yuan2024open] are both lighter than Grounded-SAM in parameter size, and therefore require less compute and memory; we report GSAM values here since it represents the most demanding case among the three. Overall, even with GSAM enabled, the complete adaptation run finishes in under $8$ hours and <18<18 GB peak memory on a single RTX A6000, confirming that the footprint remains modest.
@@
Figure A.2: Additional qualitative results. Four examples from the Foggy Cityscapes [sakaridis2018semantic] target set. Mean-feature map is obtained from taking the channel-mean from the last layer of the student’s backbone and thresholding at 0.6.Left: Baseline model  [hao2024simplifying] produces spurious background activations, leading to missed detections or localization errors (red arrows) and false positives (orange arrows). Right: Our method effectively suppresses both feature-space confusion and class-label noise, resulting in clear activations and more accurate classification and object localization. Zoom in for best view.
@@ Appendix A.6 Additional Qualitative Results
Figure A.2 presents additional qualitative examples comparing the baseline method [hao2024simplifying] and our proposed method on the Foggy Cityscapes dataset. Each row corresponds to one scene, with the first two columns illustrating the baseline’s thresholded mean-channel feature maps and predicted detections, and the last two columns showing the same representations from our method. The mean-channel map is obtained by taking the mean along the channel dimension of the last layer of the backbone and thresholding at 0.6, which is then upsampled to the image dimension for visualization. The baseline consistently exhibits dispersed activations, causing inaccuracies such as false positives (bicycle in the first image) and missed detections (car in the first image, incomplete train localization in the second, and missed buses in the third and fourth images). In contrast, our method significantly reduces the irrelevant background activations, producing cleaner, focused activations that accurately highlight relevant objects and recover detections missed by the baseline. Overall, these qualitative results illustrate that our method addresses both the pseudo-label noise and feature-space confusion that occurs due to the domain shift, resulting in improved detection accuracy and reliability.

### v3

@@ 5.2 Ablation Studies
GSAM [24] (arXiv’24) 46.2 57.2 48.4
ESC-Net [14] (CVPR’25) 46.5 57.8 49.3
OVSAM [36] (ECCV’24) 46.9 58.8 50.1
@@
Ablation on different components of IRPL. Table 7 presents an ablation study highlighting the impact of different components in IRPL (Equation 5). The peak-adjust transform ($p^{\prime}_{k}$[32] provides moderate improvement when directly applied to object detection, yielding 55.9 AP. However, better performance is achieved when it is complemented with: i) foreground-background weighting ($\mathbf{w_{\hat{c}}}$), which addresses the inherent imbalance between object and background regions and ii) entropy regularization ($\mathbf{D_{\mathrm{KL}}}$), which mitigates head-class dominance. Together, these components lead to the highest AP of 56.8, demonstrating the effectiveness of IRPL in SFOD.
@@
Table 8: Per-category AP50 results on the target-domain. We report baseline [7] performance, performance with our method, and the improvement ($\Delta$).
@@ Appendix A.2 Architecture-generality of FALCON-SFOD
To assess the architectural generality of our framework, we integrate FALCON-SFOD with multiple representative detection methods, including both conventional Faster R-CNN-based pipelines [30, 20] and transformer-based detectors such as Deformable DETR [11]. As shown in Table A.1, incorporating our proposed SPAR and IRPL modules consistently improves performance across architectures and domain-shift settings. Notably, the observed gains are obtained without modifying the detector structure or adding inference-time cost, underscoring the plug-and-play nature of our design. These results demonstrate that FALCON-SFOD generalizes well beyond a single detector family, enhancing adaptation robustness across both convolutional and transformer architectures.
@@ Appendix A.3 Reproducibility and Implementation Details
As described in Section A.2, FALCON-SFOD can be seamlessly integrated into a variety of detector architectures, consistently yielding performance gains across settings. For a fair comparison, we preserve all training parameters such as batch size, number of epochs, learning rate, optimization schedule, and data preprocessing pipeline from the corresponding baseline models, applying FALCON-SFOD as an additional adaptation module without altering the underlying training configuration. For the newly introduced hyperparameters, Section A.3.1 demonstrates that a single configuration performs consistently across benchmarks, indicating that FALCON-SFOD exhibits strong robustness and low sensitivity to hyperparameter choices.
@@ A.3.1 Hyperparameter Analysis
Imbalance-aware Noise-Robust Pseudo-Labeling. IRPL addresses the dual challenges of label noise and foreground-background imbalance inherent in source-free object detection. Unlike the standard cross-entropy objective, which can be dominated by mislabeled high-confidence predictions, IRPL utilizes a peak-adjust transform that moderates the student’s logits by adding a large margin $m$ to the most confident class and renormalizing, thereby dampening gradients for easy, clean samples while preserving full corrective signals for uncertain or mislabeled ones. This mechanism provides intrinsic robustness to noisy pseudo-labels. To further stabilize learning under imbalance, IRPL combines this with foreground-background weighting ($w_{\mathrm{fg}},w_{\mathrm{bg}}$) and a mild entropy regularization term weighted by $\gamma$, which discourages over-confident head-class predictions. Following the theoretical insights from [32], we treat the coefficients $\alpha$ and $\beta$ in
@@
as scaling factors rather than critical hyperparameters. Lemma 3 in [32] proves that the excess-risk bound of $\alpha L_{\epsilon}+\beta L_{\text{symmetric}}$ is identical to that of $\alpha L_{\epsilon}$ alone, implying that $\alpha$ and $\beta$ affect only the gradient magnitude rather than the theoretical robustness. Hence, we fix them following [32] instead of tuning.
@@ A.3.2 Results across Random Seeds
Discussion. While existing SFOD works largely report single-run results, Table A.3 provides performance across three random seeds to evaluate robustness and stability. When integrated into Simple-SFOD [7], our FALCON-SFOD consistently improves performance on the challenging $\rightarrowFoggy Cityscapes$ domain shift. Across seeds 42, 123, and 9999, our approach achieves an average mAP of 47.0 $\pm$ 0.29, demonstrating both a clear improvement over the Simple-SFOD baseline (45.0 mAP) and strong reproducibility. The small variance across runs indicates that FALCON-SFOD’s gains are stable and not sensitive to random initialization.
@@
Simple-SFOD [7] (ECCV’24) 40.9 48.0 58.9 29.6 51.9 50.2 36.2 44.1 45.0
@@ Appendix A.4 Proofs for the Lemmas and Theorems
Inspired from [32], we can write
@@ Appendix A.5 Run time and Memory Analysis
Tables A.4 and A.5 report the cost of integrating GSAM [24] into our pipeline. The time overhead is marginal: the offline GSAM pass completes in 1 0501\,050 s ($\sim$17 min), which is only $3.8\%$ of the 28 00028\,000 s required for our model training (and only $5.2\%$ relative to the baseline). When added to the full pipeline, the end-to-end wall-clock increases only slightly (28 08428\,084 s $\rightarrow$ 29 13429\,134 s), well within the typical run-to-run variability of large-scale training. The memory overhead is also short-lived: GSAM peaks at $18$ GB only during its 17-minute preprocessing, while training itself never exceeds $9.6$ GB. Since these stages do not overlap, the entire procedure fits comfortably on a single 2424-$48$ GB GPU without any modification to training. Moreover, the amortization cost is small: GSAM masks are generated once per target split and can be cached for reuse in all subsequent experiments. The extraction step is fully parallel across images, so on a multi-GPU node the elapsed time approaches standard data-loading latency. For completeness, we note that ESC-Net [14] and OV-SAM [36] are both lighter than Grounded-SAM in parameter size, and therefore require less compute and memory; we report GSAM values here since it represents the most demanding case among the three. Overall, even with GSAM enabled, the complete adaptation run finishes in under $8$ hours and <18<18 GB peak memory on a single RTX A6000, confirming that the footprint remains modest.
@@
Figure A.2: Additional qualitative results. Four examples from the Foggy Cityscapes [26] target set. Mean-feature map is obtained from taking the channel-mean from the last layer of the student’s backbone and thresholding at 0.6.Left: Baseline model  [7] produces spurious background activations, leading to missed detections or localization errors (red arrows) and false positives (orange arrows). Right: Our method effectively suppresses both feature-space confusion and class-label noise, resulting in clear activations and more accurate classification and object localization. Zoom in for best view.
@@ Appendix A.6 Additional Qualitative Results
Figure A.2 presents additional qualitative examples comparing the baseline method [7] and our proposed method on the Foggy Cityscapes dataset. Each row corresponds to one scene, with the first two columns illustrating the baseline’s thresholded mean-channel feature maps and predicted detections, and the last two columns showing the same representations from our method. The mean-channel map is obtained by taking the mean along the channel dimension of the last layer of the backbone and thresholding at 0.6, which is then upsampled to the image dimension for visualization. The baseline consistently exhibits dispersed activations, causing inaccuracies such as false positives (bicycle in the first image) and missed detections (car in the first image, incomplete train localization in the second, and missed buses in the third and fourth images). In contrast, our method significantly reduces the irrelevant background activations, producing cleaner, focused activations that accurately highlight relevant objects and recover detections missed by the baseline. Overall, these qualitative results illustrate that our method addresses both the pseudo-label noise and feature-space confusion that occurs due to the domain shift, resulting in improved detection accuracy and reliability.